15. The diameter of a round pipe cross-section 1 is 2.5 times smaller than that of cross-section 2. What would be the mean velocities vi and va? (1) Vi 1V2 (2) Vi-2V2 (3) Vi # 16V2 (4) Vi-6.25V2
Answer #1

by the continuity equation we have

A1v1 = A2v2

pi/4*d1^2 *v1 = pi/4*d2^2*v2

d1^2 * v1 = (2.5d1)^2*v2

v1 = 6.25v2

so option(4) is the correct answer